OpenStudy (love_to_love_you):

Write the equation of the line that passes through (-3, 8) and (4, -2) in a) Slope intercept form b) Point Slope form c) Standard Form Show all steps.

OpenStudy (love_to_love_you):

@iGreen

OpenStudy (anonymous):

Slope Intercept form- y=mx+b

OpenStudy (anonymous):

Point slope form: y-y1=m(x-x1) standard from- Ax+By=C

OpenStudy (love_to_love_you):

Slope inter- y=mx+b mx=slope right? b= y intercept?

OpenStudy (anonymous):

Yes, m is slope and b is the y intercept.

OpenStudy (anonymous):

So find your slope. M=y2-y1/x2-y1

OpenStudy (love_to_love_you):

Okay, I got it from here. Thank you!!!